OpenStudy (anonymous):

One solution of the equation 4x2 - 9x + 5 = 0 is five over four. What is the other solution?

OpenStudy (anonymous):

Simplifying 4x2 + 9x + 5 = 0 Reorder the terms: 5 + 9x + 4x2 = 0 Solving 5 + 9x + 4x2 = 0 Solving for variable 'x'. Factor a trinomial. (5 + 4x)(1 + x) = 0 Subproblem 1 Set the factor '(5 + 4x)' equal to zero and attempt to solve: Simplifying 5 + 4x = 0 Solving 5 + 4x = 0 Move all terms containing x to the left, all other terms to the right. Add '-5' to each side of the equation. 5 + -5 + 4x = 0 + -5 Combine like terms: 5 + -5 = 0 0 + 4x = 0 + -5 4x = 0 + -5 Combine like terms: 0 + -5 = -5 4x = -5 Divide each side by '4'. x = -1.25 Simplifying x = -1.25 Subproblem 2 Set the factor '(1 + x)' equal to zero and attempt to solve: Simplifying 1 + x = 0 Solving 1 + x = 0 Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + x = 0 + -1 Combine like terms: 1 + -1 = 0 0 + x = 0 + -1 x = 0 + -1 Combine like terms: 0 + -1 = -1 x = -1 Simplifying x = -1 Solution x = {-1.25, -1}

OpenStudy (anonymous):

So its -1 or x = {-1.25, -1}

OpenStudy (anonymous):

x can be -1 or 1.25

OpenStudy (anonymous):

ok thanx

OpenStudy (anonymous):

Welcome :)